# SpotV2Net: Multivariate Intraday Spot Volatility Forecasting via Vol-of-Vol-Informed Graph Attention Networks

### v3

@@ 2 Related literature
GNNs are part of the ML-based approaches that can play a significant role in enhancing financial applications, see, e.g., the review by Wang et al., (2021) and Zhang et al., 2022b . Recent studies (see Chen and Robert, (2022); Reisenhofer et al., (2022); Zhang et al., 2022a ; Djanga et al., (2023)) employ GNN-based models to forecast integrated measures of asset volatility, illustrating the potential of graph-based methods in capturing complex inter-asset relationships, including non-linear volatility spillover effects (Zhang et al., 2023a, ). Similarly, Wu et al., 2020b and Cheng et al., (2022) extend the scope of application for GNNs by employing the latter for forecasting financial time series. Building on these foundations, our research aims to extend the use of GNNs by including the attention mechanism typical of GATs. GAT architectures have already been employed to predict stock market movements (Kim et al.,, 2019; Cheng et al.,, 2022) and provide stock recommendations (Ying et al.,, 2020). However, few works have leveraged the flexibility of GATs for modeling spillover effects in financial markets. For instance, Cheng and Li, (2021) models the momentum effect between firms using an attribute-sensitive architecture similar to our approach, but they focus on solving a binary classification problem to predict the directional movements of stock prices, rather than forecasting asset volatilities.
@@ 5 The SpotV2Net Model
The general idea of GNNs rests on the principle of iterative message passing, where graph nodes aggregate information from themselves and their connected nodes, referred to as neighbors. This aggregation can be repeated by stacking multiple GAT layers, where each layer generates its own representation based on the output of the previous layer. These representations are then passed to a linear aggregation layer to perform node regression, i.e., obtaining output forecasts for each node. This end-to-end learning process allows training the model to operate directly on graph-structured data.
@@
$\alpha_{ij}=\text{softmax}_{j}(e_{ij})=\frac{\exp(e_{ij})}{\sum_{h=1}^{N}\exp(e_{ih})}.$
@@
$\alpha_{ij}=\frac{\exp(\mathcal{L}(q^{T}[Wx_{i}\parallel Wx_{j}]))}{\sum_{h=1}^{N}\exp(\mathcal{L}(q^{T}[Wx_{i}\parallel Wx_{h}]))},$
@@
To improve the expressiveness  [^4] Expressiveness refers to the model’s ability to capture complex patterns and relationships in the data, which is also referred to as the model’s capacity. For more details, see Chapter 5 of Goodfellow et al., (2016). and accuracy of the training process (Vaswani et al.,, 2017), when implementing SpotV2Net, we follow a multi-head attention mechanism, which consists of a number $K$ of independent attention mechanisms. Multiple heads perform the same operations on the same input but are initialized with different weights. This approach allows each head to learn distinct patterns from the inputs and capture various aspects of the data. The resulting hidden node representation is obtained via concatenation and reads as
@@
Our application of the GAT layer also allows for edge features, i.e., information regarding the dependence between two nodes in a graph. Specifically, we have a set of edge features $x^{e}=\{x_{ij}^{e}\mid i\geq 1,j\leq N,i\neq j\},\quad x_{ij}^{e}\in\mathbb{R}^{E}$. We set
@@
that is, $x_{i}^{e}$ includes contemporaneous values and all the lags up to $L$ of the estimated volatility of volatility of the $i$-th asset and the estimated co-volatility of volatility between the $i$-th asset and all the other assets. Fig. 3 illustrates the general structure of SpotV2Net’s input graph in the three-dimensional case [^5] This simplified, low-dimensional representation is intended to illustrate the process clearly, avoiding the cluttered figure that would result from using higher dimensions. However, the inherent scalability of the GNN architecture allows effortlessly applying SpotV2Net to a much larger cross-section of assets. Accordingly, in the empirical application presented in Sec. 7, we employ SpotV2Net to model the spot volatility of all of the 30 constituents of the DJIA index..
@@
$\alpha^{\prime}_{ij}=\frac{\exp\left(\mathcal{L}\left(q^{\prime T}[Wx_{i}\parallel Wx_{j}\parallel Ux_{i}^{e}]\right)\right)}{\sum_{k\in N_{i}}\exp\left(\mathcal{L}\left(q^{\prime T}[Wx_{i}\parallel Wx_{k}\parallel Ux_{i}^{e}]\right)\right)},$
where $U\in\mathbb{R}^{E^{\prime}\times E}$ is the weight matrix for the linear transformation of the edge features and $E^{\prime}$ is the dimensionality of the transformed edge features.
@@ 7 Empirical study: an application to the universe of DJIA stocks
We collect the raw price data for the 30 components of DJIA index from the TAQ - Millisecond Consolidated Trades database, which we access through the Wharton Research Data Services (WRDS). The dataset covers the period from June 1st, 2020, to May 10th, 2023, corresponding to 737 days.
For the implementation of Fourier estimator, we proceed as follows. Starting from tick data, we filter the transactions to include only those executed on the New York Stock Exchange (NYSE) and exclude transactions that occurred outside the operating market hours from 9:30 a.m. to 4:00 p.m. Then, we resample the tick data to the one-second frequency, selecting the price associated with the transaction closest to each second’s onset. The resulting sample size employed in the implementation of the estimators is $n=23400$ per day. Note that we measure time in days and set T equal to 1. The cutting frequencies $N_{v_{i}}$, $M_{v_{i}}$, $S_{v_{i}}$, $L_{v_{i}}$ (for the $i$-th asset) and $N_{c}$, $M_{c}$, $S_{c}$ and $L_{c}$ (for each pair of assets) are selected using the guidance from the paper by Sanfelici and Toscano, (2024), which introduces a MATLAB library for the implementation of Fourier estimators [^8] The Fourier method is not consistent in the presence of jumps. To remove the effect of jumps from the estimation, we identify 1-second returns whose absolute value is larger than a given threshold $\vartheta_{n}$ and replace them with zeros. Letting $\vartheta_{n}=\beta\left(T/n\right)^{\alpha}$, the constants that determines the threshold are selected using numerical simulations. As a result, we find that suitable choices are $\beta=0.5$ and $\alpha=0.5$.. Univariate and multivariate spot volatility and volatility of volatility estimates are obtained on the 30-minute grid.
@@ 7.2 Single-step forecast
Validation period (MSE) HAR-Spot XGB LSTM SpotV2Net-NE SpotV2Net HAR-Spot 6.32* 40.23* 42.12* 45.18* XGB 5.50* 6.16* 25.94* LSTM 2.34* 9.08* SpotV2Net-NE 6.12* SpotV2Net
Validation period (QLIKE) HAR-Spot XGB LSTM SpotV2Net-NE SpotV2Net HAR-Spot 13.11* 21.05* 23.45* 28.14* XGB 12.32* 13.06* 20.06* LSTM 5.67* 12.43* SpotV2Net-NE 8.91* SpotV2Net
Table 3: Comparison of single-step forecasting accuracy using the Diebold-Mariano (DM) test with MSE (top) and QLIKE (bottom) losses in the validation period. Positive numbers indicate the column model outperforms the row model. Superscript * denotes the significance levels of 5%.
Test period (MSE) HAR-Spot XGB LSTM SpotV2Net-NE SpotV2Net HAR-Spot -42.69* -1.11 -0.38 6.71* XGB 37.89* 43.25* 52.12* LSTM 0.71 11.87* SpotV2Net-NE 7.23* SpotV2Net
Test period (QLIKE) HAR-Spot XGB LSTM SpotV2Net-NE SpotV2Net HAR-Spot -44.34* -3.17* -2.41* 6.57* XGB 22.98* 26.15* 32.54* LSTM 0.55 10.62* SpotV2Net-NE 7.05* SpotV2Net
Table 4: Comparison of single-step forecasting accuracy using the Diebold-Mariano (DM) test with MSE (top) and QLIKE (bottom) losses in the test period. Positive numbers indicate the column model outperforms the row model. Superscript * denotes the significance levels of 5%.
@@ 7.3 Model Interpretation
For this purpose, in this Section we employ GNNExplainer (Ying et al.,, 2019), a model-agnostic approach that provides interpretable explanations for predictions of GNN-based models. Specifically, GNNExplainer allows the identification of the subgraph [^10] A subgraph is defined as a graph consisting of a subset of the nodes and edges of the original graph. deemed the most influential for the model’s prediction of a specific node.
@@ 7.4 Multi-step forecast
To evaluate functional forecasts accuracy, we use $global$ loss functions. Specifically, we employ the average MSE of the multi-step prediction, that is,
@@
Validation period (MSE) HAR-Spot XGB LSTM SpotV2Net-NE SpotV2Net HAR-Spot 7.12* 46.76* 48.21* 58.23* XGB 28.55* 35.15* 47.36* LSTM 5.64* 11.98* SpotV2Net-NE 8.53* SpotV2Net
Validation period (QLIKE) HAR-Spot XGB LSTM SpotV2Net-NE SpotV2Net HAR-Spot 5.41* 30.20* 37.05* 43.82* XGB 24.79* 31.63* 38.37* LSTM 6.83* 14.51* SpotV2Net-NE 7.09* SpotV2Net
Table 6: Comparison of multi-step forecasting accuracy using the Diebold-Mariano (DM) test with average MSE (top) and QLIKE (bottom) losses on the validation set. Positive numbers indicate the column model outperforms the row model. Superscript * denotes the significance levels of 5%.
Test period (MSE) HAR-Spot XGB LSTM SpotV2Net-NE SpotV2Net HAR-Spot -25.57* 9.80* 13.52* 33.16* XGB 35.38* 39.10* 59.51* LSTM 3.71* 24.61* SpotV2Net-NE 21.01* SpotV2Net
Test period (QLIKE) HAR-Spot XGB LSTM SpotV2Net-NE SpotV2Net HAR-Spot -20.23* 7.90* 11.15* 28.57* XGB 28.21* 31.34* 46.95* LSTM 3.13* 21.34* SpotV2Net-NE 17.15* SpotV2Net
Table 7: Comparison of multi-step forecasting accuracy using the Diebold-Mariano (DM) test with average MSE (top) and QLIKE (bottom) losses on the validation set. Positive numbers indicate the column model outperforms the row model. Superscript * denotes the significance levels of 5%.
@@ A.3 Long Short-term Memory (LSTM) Neural Network
The LSTM architecture proposes a memory block to capture long-term dependencies more effectively. Within each LSTM unit, a standard transformation of the original input occurs.
Consider the vector of input features $\mathbf{x}_{\tau_{b}}$ at time $\tau_{b}$, which comprises all companies’ univariate volatilities and co-volatilities. The LSTM model accepts these vectors $\mathbf{x}_{\tau_{b}}$ up to a given lag $L$. Note that we input only spot volatilities and co-volatilities time series in the LSTM algorithm because our experiments suggest that also inputting univariate and multivariate co-volatilities of volatilities and their lags worsens the model’s performance. We attribute this performance decrease to the large dimensionality of the input size. In contrast, SpotV2Net uses this information efficiently by distributing it into a graph structure with node and edge features. LSTM has no knowledge of such structure and thus infers relationships by combining information through different layers, which proves to be limited in our empirical experiment.

### v4

@@ 2 Related literature
GNNs are part of the ML-based approaches that can play a significant role in enhancing financial applications, see, e.g., the review by Wang et al., (2021) and Zhang et al., 2022b . Recent studies (see Chen and Robert, (2022); Reisenhofer et al., (2022); Zhang et al., 2022a ; Djanga et al., (2023)) employ GNN-based models to forecast integrated measures of asset volatility, illustrating the potential of graph-based methods in capturing complex inter-asset relationships, including non-linear volatility spillover effects (Zhang et al., 2023a, ). Similarly, Wu et al., 2020b and Cheng et al., (2022) extend the scope of application for GNNs by employing the latter for forecasting financial time series. Building on these foundations, our research aims to extend the use of GNNs by including the attention mechanism typical of GATs. GAT architectures have already been employed to predict stock market movements (Kim et al.,, 2019; Cheng et al.,, 2022) and provide stock recommendations (Ying et al.,, 2020). However, few works have leveraged the flexibility of GATs for modeling spillover effects in financial markets. For instance, Cheng and Li, (2021) model the momentum effect between firms using an attribute-sensitive architecture similar to our approach, but they focus on solving a binary classification problem to predict the directional movements of stock prices, rather than forecasting asset volatilities.
@@ 5 The SpotV2Net Model
The general idea of GNNs rests on the principle of iterative message passing, where graph nodes aggregate information from themselves and their connected nodes, referred to as neighbors. This aggregation can be repeated by stacking multiple layers, where each layer generates its own representation based on the output of the previous layer. These representations are then passed to a linear aggregation layer to perform node regression, i.e., obtaining output forecasts for each node. This end-to-end learning process allows training the model to operate directly on graph-structured data.
@@
$\alpha_{ij}=\text{softmax}_{j}(e_{ij})=\frac{\exp(e_{ij})}{\sum_{k=1}^{N}\exp(e_{ik})}.$
@@
$\alpha_{ij}=\frac{\exp(\mathcal{L}(q^{T}[Wx_{i}\parallel Wx_{j}]))}{\sum_{k=1}^{N}\exp(\mathcal{L}(q^{T}[Wx_{i}\parallel Wx_{k}]))},$
@@
To improve the expressiveness  [^4] Expressiveness (or capacity) refers to the model’s ability to capture complex patterns and relationships in the data. For more details, see Chapter 5 of Goodfellow et al., (2016). and accuracy of the training process (Vaswani et al.,, 2017), when implementing SpotV2Net, we follow a multi-head attention mechanism, which consists of a number $K$ of independent attention mechanisms. Multiple heads perform the same operations on the same input but are initialized with different weights. This approach allows each head to learn distinct patterns from the inputs and capture various aspects of the data. The resulting hidden node representation is obtained via concatenation and reads as
@@
Our application of the GAT layer also allows for edge features, i.e., information regarding the dependence between two nodes in a graph. Specifically, we have a set of edge features $x^{e}=\{x_{ij}^{e}\mid i<j\},\quad x_{ij}^{e}\in\mathbb{R}^{E}$. We set
@@
that is, $x_{ij}^{e}$ includes contemporaneous values and all the lags up to $L$ of the estimated volatility of volatility of the $i$-th asset and the $j$-th asset, along with the estimated co-volatility of volatility between the $i$-th asset and the $j$-th asset. Fig. 3 illustrates the general structure of SpotV2Net’s input graph in the three-dimensional case [^5] This simplified, low-dimensional representation is intended to illustrate the process clearly, avoiding the cluttered figure that would result from using higher dimensions. However, the inherent scalability of the GNN architecture allows effortlessly applying SpotV2Net to a much larger cross-section of assets. Accordingly, in the empirical application presented in Sec. 7, we employ SpotV2Net to model the spot volatility of all of the 30 constituents of the DJIA index..
@@
$\alpha^{\prime}_{ij}=\frac{\exp\left(\mathcal{L}\left(q^{\prime T}[Wx_{i}\parallel Wx_{j}\parallel U{x_{ij}^{e}}]\right)\right)}{\sum_{k=1}^{N}{\exp\left(\mathcal{L}\left(q^{\prime T}[Wx_{i}\parallel Wx_{k}\parallel U{x_{ik}^{e}}]\right)\right)}},$
where $U\in\mathbb{R}^{E^{\prime}\times E}$ is the weight matrix for the linear transformation of the edge features and $E^{\prime}$ is the dimension of the transformed edge features.
@@ 7 Empirical study: an application to the universe of DJIA stocks
We collect the raw price data for the 30 components of the DJIA index from the TAQ - Millisecond Consolidated Trades database, which we access through the Wharton Research Data Services (WRDS). The dataset covers the period from June 1st, 2020, to May 10th, 2023, corresponding to 737 days.
For the implementation of Fourier estimator, we proceed as follows. Starting from tick data, we filter the transactions to include only those executed on the New York Stock Exchange (NYSE) and exclude transactions that occurred outside the operating market hours from 9:30 a.m. to 4:00 p.m. Then, we resample the tick data to the one-second frequency, selecting the price associated with the transaction closest to each second’s onset. The resulting sample size employed in the implementation of the estimators is $n=23400$ per day. Note that we measure time in days and set T equal to 1. The cutting frequencies $N_{v_{i}}$, $M_{v_{i}}$, $S_{v_{i}}$, $L_{v_{i}}$ (for the $i$-th asset) and $N_{c}$, $M_{c}$, $S_{c}$ and $L_{c}$ (for each pair of assets) are selected using the guidance from the paper by Sanfelici and Toscano, (2024), which introduces a MATLAB library for the implementation of Fourier estimators [^8] The Fourier method is not consistent in the presence of jumps. To remove the effect of jumps from the estimation, we identify 1-second returns whose absolute value is larger than a given threshold $\vartheta_{n}$ and replace them with zeros. Letting $\vartheta_{n}=\beta\left(T/n\right)^{\alpha}$, the constants that determine the threshold are selected using numerical simulations. As a result, we find that suitable choices are $\beta=0.5$ and $\alpha=0.5$.. Univariate and multivariate spot volatility and volatility of volatility estimates are obtained on the 30-minute grid.
@@ 7.2 Single-step forecast
Validation period (MSE)
XGB LSTM SpotV2Net-NE SpotV2Net
HAR-Spot 6.32* 40.23* 42.12* 45.18*
XGB 5.50* 6.16* 25.94*
LSTM 2.34* 9.08*
SpotV2Net-NE 6.12*
Validation period (QLIKE)
XGB LSTM SpotV2Net-NE SpotV2Net
HAR-Spot 13.11* 21.05* 23.45* 28.14*
XGB 12.32* 13.06* 20.06*
LSTM 5.67* 12.43*
SpotV2Net-NE 8.91*
Table 3: Comparison of single-step forecasting accuracy using the DM test with MSE (top) and QLIKE (bottom) losses in the validation period. Positive numbers indicate the column model outperforms the row model. Superscript * denotes the significance levels of 5%.
Test period (MSE)
XGB LSTM SpotV2Net-NE SpotV2Net
HAR-Spot -42.69* -1.11 -0.38 6.71*
XGB 37.89* 43.25* 52.12*
LSTM 0.71 11.87*
SpotV2Net-NE 7.23*
Test period (QLIKE)
XGB LSTM SpotV2Net-NE SpotV2Net
HAR-Spot -44.34* -3.17* -2.41* 6.57*
XGB 22.98* 26.15* 32.54*
LSTM 0.55 10.62*
SpotV2Net-NE 7.05*
Table 4: Comparison of single-step forecasting accuracy using the DM test with MSE (top) and QLIKE (bottom) losses in the test period. Positive numbers indicate the column model outperforms the row model. Superscript * denotes the significance levels of 5%.
@@ 7.3 Model interpretation
For this purpose, we employ GNNExplainer (Ying et al.,, 2019), a model-agnostic approach that provides interpretable explanations for predictions of GNN-based models. Specifically, GNNExplainer allows the identification of the subgraph [^10] A subgraph is defined as a graph consisting of a subset of the nodes and edges of the original graph. deemed the most influential for the model’s prediction of a specific node.
@@ 7.4 Multi-step forecast
To evaluate functional forecast accuracy, we use $global$ loss functions. Specifically, we employ the average MSE of the multi-step prediction, that is,
@@
Validation period (MSE)
XGB LSTM SpotV2Net-NE SpotV2Net
HAR-Spot 7.12* 46.76* 48.21* 58.23*
XGB 28.55* 35.15* 47.36*
LSTM 5.64* 11.98*
SpotV2Net-NE 8.53*
Validation period (QLIKE)
XGB LSTM SpotV2Net-NE SpotV2Net
HAR-Spot 5.41* 30.20* 37.05* 43.82*
XGB 24.79* 31.63* 38.37*
LSTM 6.83* 14.51*
SpotV2Net-NE 7.09*
Table 6: Comparison of multi-step forecasting accuracy using the DM test with average MSE (top) and QLIKE (bottom) losses on the validation set. Positive numbers indicate the column model outperforms the row model. Superscript * denotes the significance levels of 5%.
Test period (MSE)
XGB LSTM SpotV2Net-NE SpotV2Net
HAR-Spot -25.57* 9.80* 13.52* 33.16*
XGB 35.38* 39.10* 59.51*
LSTM 3.71* 24.61*
SpotV2Net-NE 21.01*
Test period (QLIKE)
XGB LSTM SpotV2Net-NE SpotV2Net
HAR-Spot -20.23* 7.90* 11.15* 28.57*
XGB 28.21* 31.34* 46.95*
LSTM 3.13* 21.34*
SpotV2Net-NE 17.15*
Table 7: Comparison of multi-step forecasting accuracy using the DM test with average MSE (top) and QLIKE (bottom) losses on the test set. Positive numbers indicate the column model outperforms the row model. Superscript * denotes the significance levels of 5%.
@@ 8 Conclusions
Acknowledgements
The authors would like to thank the Associate Editor and two anonymous Referees for their valuable comments and suggestions. Furthermore, the authors gratefully acknowledge financial support from the Institut Louis Bachelier 2022 grant ‘Risk management in times of unprecedented geo-political volatility: a machine learning approach’.
@@ A.3 Long Short-term Memory (LSTM) Neural Network
The LSTM architecture proposes a memory block to capture long-term dependencies more effectively. Within each LSTM unit, a standard transformation of the original input occurs. Consider the vector of input features $\mathbf{x}_{\tau_{b}}$ at time $\tau_{b}$, which comprises all companies’ univariate volatilities and co-volatilities. The LSTM model accepts these vectors $\mathbf{x}_{\tau_{b}}$ up to a given lag $L$. Note that we input only spot volatilities and co-volatilities time series in the LSTM algorithm because our experiments suggest that also inputting univariate and multivariate co-volatilities of volatilities and their lags worsens the model’s performance. We attribute this performance decrease to the large dimensionality of the input size. In contrast, SpotV2Net uses this information efficiently by distributing it into a graph structure with node and edge features. LSTM has no knowledge of such structure and thus infers relationships by combining information through different layers, which proves to be limited in our empirical experiment.